A Complete Bibliography of The American Statistician: 1940–1949

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

27 July 2020 Version 1.01

Title word cross-reference

107th [Ano47a, Ano47b]. 108th [Ano48a, Ano48b]. 109th [Ano49a]. 1948 [DN48].

A. [Ano47c]. Accidents [HJ48]. According [Kel49]. Action [Guy48]. Activities [Ano49i]. Affairs [Ano48-33]. Affirmative [Eke49]. against [Mye47]. Agriculture [Eck48, Jen48]. American [Ano47b, Ano48a, Ano48b, Ano49a, Wil47b]. Analysis [Ano49y, Eps47]. Annual [Ano47a, Ano47b, Ano48a, Ano48b, Ano49a, Ano49p]. Answer [MF48, Tuk49a]. Answers [Mos47, MF47, MWF48, MH48, MBM^{+}48, MGG48, MF48, MA48, MWK49, ME49, Mos49, MT49, MMH49, PTM47]. Area [Hae49a]. ASA [Ano49i]. Association [Ano48h, Ano47b, Ano48a, Ano48b, Ano49a, Ano49g, Ste48a, Wil47b]. Attitudes [SM47]. Audience [Kel49]. Authority [Cop48]. Averages [Ano49h, Arm49].

Back [Ano47d, Ano47e]. Bar [Hae48b]. Basis [Hor48]. Bias [Hae49a]. Bill [Ano47k]. Blackout [Ano48-31]. Book [Kyd48]. Boss [Ste48b]. Bradstreet [Hay48]. Bureau [Ded48, Ano48j]. Business [Ano49p].

Cards [Eps47]. Cartographic [Hit48]. Case [Eke49, Mye47]. Casual [Mye47]. Census [DN48, Ded48]. Censuses [Eck48]. Chapter [Ano47f, Ano48c, Ano48d, Ano48e, Ano48f, Ano48g, Ano49b, Ano49c, Ano49d, Ano49e, Ano49i, Hoa47, Wat47]. Chapters [Tuk49b, Tuk49c]. Chart [Hae47a, Hae49b, Kel49]. Charts [Hae48b, Hae49b, Heg48, Mye47]. Chicago [Ano49i]. China [Lie48]. Classification [Dem48, Hay48]. Classified [Kel49]. Collecting [Lie48]. College [Dod48]. Combining [MF48]. Commission [Ano48-36]. Committee [Ano49f, Ano49y, Ano49j]. Committees [Ano48h, Ano49g, Ano49f, Ano49j]. Common [Kel49]. Comparisons [Wie48]. Computing [Alt47]. Conference [Sha48]. Conferences [Har47]. Consultant [Eis48]. Consumer [SM47]. Control [HJ48]. Coordination [Cop48, Lin49a, Lin49b]. Correction [Ano49h]. Correlation [Uri48]. Council [Ano49y]. Criminal [Wie48]. Crisis [Wil47a]. Current [Ano47g]. Curriculum [NA47].

Dangerous [Hae48a]. Decisions [Hor48]. Department [Jen48]. Design [Sne48]. Development [Lin49a, Lin49b]. Devices [Alt47, Kel49]. Discovered [Ken48a]. Display [Heg48]. Does [Ano48i]. Don’t [Mye47]. Doomsday [Kyd48]. Double [Hae48a]. Dun [Hay48].

Ease [Kel49]. East [DN48]. Economic [Leh48]. Editor [Cal49, YA49]. Education [Guy48]. efficiency [Tuk49a]. Election [Ano49y]. Employment [Giv49, Ano48j]. Engineering [Ano48-34, Tum47]. England [Soc48]. estimates [Tuk49a]. Everyday [Ano48-33]. Examination [Ken48b]. Executive [Hor48]. Expanding [Ano49i]. Extend [Eke49]. Eyes [Mye47].

Factor [Hit48]. FAO [Tae48]. Federal [Giv49, Ano47g, Ano48-32, Lin49a, Lin49b, Tuk49b, Tuk49c]. Field [RCV48]. First [Ano49j]. Fiscal [Ano47g]. Forecasts [Ano49y]. Foreign [Ded48]. Foundation [Ano47k]. Founder [Wil47b]. Front [Ano47h, Ano47i, Ano47j, Ano48k, Ano48l, Ano48m, Ano48n, Ano48o, Ano48p, Ano49k, Ano49l, Ano49m, Ano49n, Ano49o].

Gaps [Leh48]. Government [Tuk49b, Tuk49c]. Graduate [Jen48]. Graphic [Kel49, Uri48].

Hand [Eps47]. Hemisphere [Guy48]. High [Alt47]. High-Speed [Alt47]. Highlights [Ano49p]. Hold [Hae47a]. Hospitals [Den47].

If [Ste48b]. II [Tuk49c]. Illustrative [Giv49]. Implications [Ano48-30].

Independent [MF48]. Indicators [Leh48]. Industrial [Hay48]. Infinite [Tuk47]. Information [Sha48]. Institute [Ric49]. Instruments [Cop48]. International [Har47, Ric49]. Internships [Jen48]. Interviewers [RCV48]. Investigations [Sne48]. Iowa [Dod48]. IV [Tuk49b].

Labor [Ano48j, Giv49]. Latin [Ken48a]. Laymen [Mye47]. Lemuel [Wil47b]. Letter [Cal49, YA49]. Libraries [Ste48a]. Life [Ano48i]. likelihood [Tuk49a]. Line [Hae47a]. Linear [Uri48]. Log [Mye47]. Logarithmic [Mye47]. Lotteries [Bla49].

Management [Sea48]. Map [Hae49a]. Market [Giv49]. Marketing [RCV48]. Matter [Ano47d, Ano47e, Ano47h, Ano47i, Ano47j, Ano48k, Ano48l, Ano48m, Ano48n, Ano48o, Ano48p, Ano49k, Ano49l, Ano49m, Ano49n, Ano49o]. Maximum [Tuk49a]. Meeting [Ano47b, Ano48a, Ano48b, Ano49a, Ano47a, Ano49p]. Members [Ano47o, Ano47p, Ano47q, Ano48x, Ano48y, Ano48z, Ano48-27, Ano48-28, Ano49v, Ano49w, Ano49x]. Membership [Lea49]. Memorandum [Tuk49b, Tuk49c]. Modern [Sea48]. Moving [Ano49h, Arm49]. Multiple [Uri48].

National [Ano47k]. Nations [Ano48-35, Ano48-36]. News [Ano47o, Ano47p, Ano47q, Ano47l, Ano47m, Ano47n, Ano48x, Ano48y, Ano48z, Ano48-27, Ano48-28, Ano48r, Ano48s, Ano48t, Ano48u, Ano48v, Ano48w, Ano48-35, Ano49v, Ano49w, Ano49x, Ano49q, Ano49r, Ano49s, Ano49t, Ano49u]. Notes [Ano47f, Ano48c, Ano48d, Ano48e, Ano48f, Ano48g, Ano49b, Ano49c, Ano49d, Ano49e, Giv49, Hoa47, Wat47].

observations [Tuk47]. One [Ano48-29, Hae49b]. Operations [Ano48-30, Hor48]. Opinion [SM47]. Organization [Eis48]. Outlook [Zap48]. Overlapping [Lea49].

Part [Tuk49b, Tuk49c]. Peacetime [Ano48-30]. Perils [Hae47b]. Personnel [Hat47, Wil47a]. Perspective [Hae47b, Ste48c, Ste49a, Ste49b]. Physical [Ano48-34]. Pictorial [Jah48]. Placement [Hat47, Ste48a]. Planning [DN48, Eck48]. Plea [Hae47a]. Polls [Ano49y]. Pools [Bla49]. Population [DN48, Eck48]. Positions [Ano48-32]. Practices [Den47]. Pre [Ano49y]. Pre-Election [Ano49y]. Presentation [Hae49a]. Preservation [Hae47a]. Presidency [Ano48i]. Price [Hae49b]. Problems [DN48]. Professional [Hat47]. Program [Ded48, Dod48, Guy48, Tae48, Ano48j]. Programs [Ano47g, Ano47c]. Promoting [Guy48]. Punched [Eps47].

Question [MF48, Tuk47, Tuk49a]. Questions [Mos47, MF47, MWF48, MH48, MBM^{+}48, MGG48, MF48, MA48, MWK49, ME49, Mos49, MT49, MMH49, PTM47].

Range [Hae48b]. Range-Bar [Hae48b]. Reading [Kel49]. Reason [Cop48]. Reference [Giv49]. Relations [Giv49]. Report [Ano49j]. Representation [Hit48, Uri48]. Representatives [Ano48h, Ano49g]. Research [Ano48-30, Ano49y, Eis48, Hor48, RCV48]. Role [Eis48]. Rounded [NA47]. Royal [Ken48b, Sha48]. Rulings [Hae47a].

S [Ano47c]. Sampling [Jen48, Sne48, SM47, Ano48-36]. Scale [Hae47a, Hae49b]. Scales [Eke49, Hae48a]. School [Jen48]. Science [Ano47k, Ano48-29, Ano49y]. Sciences [Ano48-34]. Scientific [Hor48, Sha48]. Scotland [Kyd48]. Section [Ano48q]. Selection [RCV48]. Semi [Mye47]. Service [Ste48a, Ano48-32]. Set [Heg48]. Shattuck [Wil47b]. Shorten [Ano48i]. Should [Eke49]. Significance [MF48]. Social [Ano49y]. Societies [Ano47r, Hat47, Lea49]. Society [Sha48, Ken48b]. Some [DN48]. Sorted [Eps47]. Special [Ste48a]. Specialized [Ano47r]. Speed [Alt47]. Sponsors [Jen48]. Square [Ken48a]. State [Dod48, Giv49]. States [Cop48]. Statist [Wil47b]. Statistical [Ano47b, Ano47g, Ano47r, Ano48a, Ano48b, Ano48-31, Ano48-32, Ano48-35, Ano48-36, Ano49a, Den47, Dod48, Eis48, Eps47, Guy48, Har47, HJ48, Ken48b, Leh48, Lin49a, Lin49b, Ric49, Tae48, Tum47, Wil47a, Wil47b]. Statistician [Ano48-33, Sea48]. Statisticians [Ano47s, Dem48, Ded48]. Statistics [Ano48j, Ano48q, Ano48-34, Giv49, Lie48, NA47, Soc48, Tuk49b, Tuk49c, Zap48, Ano48j]. Sub [Ano48-36]. Sub-Commission [Ano48-36]. Suiting [Kel49]. Supervision [RCV48]. Supplementary [Hae49b]. Supplementary-Scale [Hae49b]. Symbols [Jah48].

Teaching [Soc48]. Tests [MF48]. Three [Hit48]. Three-Factor [Hit48]. Training [Ano47s, Ano48q, Ded48, RCV48, Tae48]. Two [Hae49b].

U.S. [Jen48]. Unemployment [Giv49]. United [Ano48-35, Ano48-36, Cop48]. Use [Heg48, Jah48, Mye47].

V [Tuk49c]. VII [Tuk49c].

Wants [SM47]. Welcome [Har47]. Well [NA47]. Well-Rounded [NA47]. Were [Ste48b]. Western [Guy48]. Who [Ken48a]. Women [Zap48]. World [Ano48-29].

Year [Ano47g]. York [Ste48c, Ste49a, Ste49b].

Zero [Eke49].

References

Alt:1947:NHS

[Alt47] Franz L. Alt. New high-speed computing devices. The American Statistician, 1(1):14–15, August 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681523.

Anonymous:1947:AM

[Ano47a] Anonymous. 107th Annual Meeting. The American Statistician, 1(2):5, October 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681621.

Anonymous:1947:AMA

[Ano47b] Anonymous. 107th Annual Meeting of the American Statistical Association. The American Statistician, 1(3):4–5, December 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682531.

Anonymous:1947:P

[Ano47c] Anonymous. A. S. A. programs. The American Statistician, 1(1): 5–6, August 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681520.

Anonymous:1947:BMa

[Ano47d] Anonymous. Back matter. The American Statistician, 1(1):??, August 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681529.

Anonymous:1947:BMb

[Ano47e] Anonymous. Back matter. The American Statistician, 1(2):??, October 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681629.

Anonymous:1947:CN

[Ano47f] Anonymous. Chapter notes. The American Statistician, 1(3):21–22, December 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682539.

Anonymous:1947:FSP

[Ano47g] Anonymous. Federal statistical programs for the current fiscal year. The American Statistician, 1(3):13–16, December 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682535.

Anonymous:1947:FMa

[Ano47h] Anonymous. Front matter. The American Statistician, 1(1):1, August 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681517.

Anonymous:1947:FMb

[Ano47i] Anonymous. Front matter. The American Statistician, 1(2):??, October 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681618.

Anonymous:1947:FMc

[Ano47j] Anonymous. Front matter. The American Statistician, 1(3):??, December 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682528.

Anonymous:1947:NSF

[Ano47k] Anonymous. The National Science Foundation bill. The American Statistician, 1(2):1, October 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681619.

Anonymous:1947:Na

[Ano47l] Anonymous. News. The American Statistician, 1(1):2–4, August 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681518.

Anonymous:1947:Nb

[Ano47m] Anonymous. News. The American Statistician, 1(2):2–4, October 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681620.

Anonymous:1947:Nc

[Ano47n] Anonymous. News. The American Statistician, 1(3):2–9, December 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682530.

Anonymous:1947:NAMa

[Ano47o] Anonymous. News about members. The American Statistician, 1(1):23–24, August 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681527.

Anonymous:1947:NAMb

[Ano47p] Anonymous. News about members. The American Statistician, 1(2):15–16, October 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681627.

Anonymous:1947:NAMc

[Ano47q] Anonymous. News about members. The American Statistician, 1(3):20, December 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682538.

Anonymous:1947:SSS

[Ano47r] Anonymous. Specialized statistical societies. The American Statistician, 1(3):1, December 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682529.

Anonymous:1947:TS

[Ano47s] Anonymous. The training of statisticians. The American Statistician, 1(3):8–9, December 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682533.

Anonymous:1948:AMAa

[Ano48a] Anonymous. 108th Annual Meeting of the American Statistical Association. The American Statistician, 2(3):4–6, June 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682081.

Anonymous:1948:AMAb

[Ano48b] Anonymous. 108th Annual Meeting of the American Statistical Association. The American Statistician, 2(5):20–28, October 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681649.

Anonymous:1948:CNa

[Ano48c] Anonymous. Chapter notes. The American Statistician, 2(2):25–26, April 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682236.

Anonymous:1948:CNb

[Ano48d] Anonymous. Chapter notes. The American Statistician, 2(3):26, June 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682090.

Anonymous:1948:CNc

[Ano48e] Anonymous. Chapter notes. The American Statistician, 2(4):26, August 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682692.

Anonymous:1948:CNd

[Ano48f] Anonymous. Chapter notes. The American Statistician, 2(5):34, October 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681653.

Anonymous:1948:CNe

[Ano48g] Anonymous. Chapter notes. The American Statistician, 2(6):22, December 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681562.

Anonymous:1948:CAR

[Ano48h] Anonymous. Committees and association representatives for 1948. The American Statistician, 2(3):21, June 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682086.

Anonymous:1948:DPS

[Ano48i] Anonymous. Does the Presidency shorten life? The American Statistician, 2(2):14–15, April 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682230.

Anonymous:1948:ESP

[Ano48j] Anonymous. The Employment Statistics Program of the Bureau of Labor Statistics. The American Statistician, 2(2):21–22, April 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682233.

Anonymous:1948:FMa

[Ano48k] Anonymous. Front matter. The American Statistician, 2(1):??, February 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682357.

Anonymous:1948:FMb

[Ano48l] Anonymous. Front matter. The American Statistician, 2(2):20, April 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682224.

Anonymous:1948:FMc

[Ano48m] Anonymous. Front matter. The American Statistician, 2(3):??, June 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682078.

Anonymous:1948:FMd

[Ano48n] Anonymous. Front matter. The American Statistician, 2(4):??, August 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682677.

Anonymous:1948:FMe

[Ano48o] Anonymous. Front matter. The American Statistician, 2(5):??, October 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681641.

Anonymous:1948:FMf

[Ano48p] Anonymous. Front matter. The American Statistician, 2(6):??, December 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681551.

Anonymous:1948:NST

[Ano48q] Anonymous. A new section on training in statistics. The American Statistician, 2(1):1, February 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682358.

Anonymous:1948:Na

[Ano48r] Anonymous. News. The American Statistician, 2(1):2–22, February 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682359.

Anonymous:1948:Nb

[Ano48s] Anonymous. News. The American Statistician, 2(2):2–5, April 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682226.

Anonymous:1948:Nc

[Ano48t] Anonymous. News. The American Statistician, 2(3):2–23, June 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682080.

Anonymous:1948:Nd

[Ano48u] Anonymous. News. The American Statistician, 2(4):2–4, August 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682679.

Anonymous:1948:Ne

[Ano48v] Anonymous. News. The American Statistician, 2(5):2–29, October 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681643.

Anonymous:1948:Nf

[Ano48w] Anonymous. News. The American Statistician, 2(6):2–5, December 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681553.

Anonymous:1948:NAMa

[Ano48x] Anonymous. News about members. The American Statistician, 2(2):24–25, April 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682235.

Anonymous:1948:NAMb

[Ano48y] Anonymous. News about members. The American Statistician, 2 (3):25, June 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682089.

Anonymous:1948:NAMc

[Ano48z] Anonymous. News about members. The American Statistician, 2(4):24, August 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682690.

Anonymous:1948:NAMd

[Ano48-27] Anonymous. News about members. The American Statistician, 2(5):33–34, October 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681652.

Anonymous:1948:NAMe

[Ano48-28] Anonymous. News about members. The American Statistician, 2(6):20–21, December 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681560.

Anonymous:1948:OWS

[Ano48-29] Anonymous. One world of science. The American Statistician, 2(4): 6–7, August 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682681.

Anonymous:1948:PIO

[Ano48-30] Anonymous. Peacetime implications of operations research. The American Statistician, 2(6):1–18, December 1948. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681552.

Anonymous:1948:SB

[Ano48-31] Anonymous. Statistical blackout. The American Statistician, 2(3): 1, June 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682079.

Anonymous:1948:SPF

[Ano48-32] Anonymous. Statistical positions in the Federal Service. The American Statistician, 2(2):1–13, April 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682225.

Anonymous:1948:SEA

[Ano48-33] Anonymous. The statistician and everyday affairs. The American Statistician, 2(5):1, October 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681642.

Anonymous:1948:SPS

[Ano48-34] Anonymous. Statistics, the physical sciences and engineering. The American Statistician, 2(4):1, August 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682678.

Anonymous:1948:UNSa

[Ano48-35] Anonymous. United nations statistical news. The American Statistician, 2(4):5–6, August 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682680.

Anonymous:1948:UNSb

[Ano48-36] Anonymous. United Nations Sub-Commission on Statistical Sampling. The American Statistician, 2(4):10–19, August 1948. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682683.

Anonymous:1949:AMA

[Ano49a] Anonymous. 109th Annual Meeting of the American Statistical Association. The American Statistician, 3(4):4–9, October 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681355.

Anonymous:1949:CNa

[Ano49b] Anonymous. Chapter notes. The American Statistician, 3(1):26, February 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681397.

Anonymous:1949:CNb

[Ano49c] Anonymous. Chapter notes. The American Statistician, 3(2):21–22, April/May 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682743.

Anonymous:1949:CNc

[Ano49d] Anonymous. Chapter notes. The American Statistician, 3(3):14, June/July 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682142.

Anonymous:1949:CNd

[Ano49e] Anonymous. Chapter notes. The American Statistician, 3(5):13–14, December 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682411.

Anonymous:1949:CC

[Ano49f] Anonymous. The committee on committees. The American Statistician, 3(2):1, April/May 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682734.

Anonymous:1949:CAR

[Ano49g] Anonymous. Committees and Association representatives for 1949. The American Statistician, 3(2):17, April/May 1949. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682740.

Anonymous:1949:CMA

[Ano49h] Anonymous. Correction: Moving averages. The American Statistician, 3(5):4, December 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682406.

Anonymous:1949:EAC

[Ano49i] Anonymous. Expanding activities of the Chicago Chapter of the ASA. The American Statistician, 3(4):13–14, October 1949. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681358.

Anonymous:1949:FRC

[Ano49j] Anonymous. First report of the Committee on Committees. The American Statistician, 3(5):1–9, December 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682404.

Anonymous:1949:FMa

[Ano49k] Anonymous. Front matter. The American Statistician, 3(1):??, February 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681387.

Anonymous:1949:FMb

[Ano49l] Anonymous. Front matter. The American Statistician, 3(2):??, April/May 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682733.

Anonymous:1949:FMc

[Ano49m] Anonymous. Front matter. The American Statistician, 3(3):??, June/July 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682135.

Anonymous:1949:FMd

[Ano49n] Anonymous. Front matter. The American Statistician, 3(4):??, October 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681352.

Anonymous:1949:FMe

[Ano49o] Anonymous. Front matter. The American Statistician, 3(5):??, December 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682403.

Anonymous:1949:HAB

[Ano49p] Anonymous. Highlights of the Annual Business Meeting. The American Statistician, 3(1):1, February 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681388.

Anonymous:1949:Na

[Ano49q] Anonymous. News. The American Statistician, 3(1):2–21, February 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681389.

Anonymous:1949:Nb

[Ano49r] Anonymous. News. The American Statistician, 3(2):2–20, April/ May 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682735.

Anonymous:1949:Nc

[Ano49s] Anonymous. News. The American Statistician, 3(3):2–5, June/ July 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682137.

Anonymous:1949:Nd

[Ano49t] Anonymous. News. The American Statistician, 3(4):2–11, October 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681354.

Anonymous:1949:Ne

[Ano49u] Anonymous. News. The American Statistician, 3(5):2–5, December 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682405.

Anonymous:1949:NAMa

[Ano49v] Anonymous. News about members. The American Statistician, 3(1):25–26, February 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681396.

Anonymous:1949:NAMb

[Ano49w] Anonymous. News about members. The American Statistician, 3(2):22, April/May 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682744.

Anonymous:1949:NAMc

[Ano49x] Anonymous. News about members. The American Statistician, 3(5):12–13, December 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682410.

Anonymous:1949:SSR

[Ano49y] Anonymous. Social Science Research Council Committee on Analysis of Pre-Election Polls and Forecasts. The American Statistician, 3(1):4–5, February 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681390.

Armstrong:1949:MA

[Arm49] Charles E. Armstrong. Moving averages. The American Statistician, 3(4):10, October 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681356.

Blanch:1949:LP

[Bla49] Ernest Blanch. Lotteries and pools. The American Statistician, 3(1):18–21, February 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681392.

Callander:1949:LE

[Cal49] W. F. Callander. Letter to the Editor. The American Statistician, 3(4):1, October 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681353.

Copeland:1948:ARI

[Cop48] Morris A. Copeland. Authority and reason as instruments of coordination in the United States. The American Statistician, 2(3):7–20, June 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682082.

Dedrick:1948:CBT

[Ded48] Calvert L. Dedrick. The Census Bureau Training Program for Foreign Statisticians. The American Statistician, 2(5):13–15, October 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681647.

Deming:1948:CS

[Dem48] W. Edwards Deming. On the classification of statisticians. The American Statistician, 2(2):16–19, April 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682231.

Densen:1947:SPH

[Den47] Paul M. Densen. Statistical practices in hospitals. The American Statistician, 1(3):10–16, December 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682534.

Department:1948:EAP

[DN48] East African Statistical Department and Nairobi. East African Population Census 1948: Some problems of planning. The American Statistician, 2(4):17–18, August 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682686.

Dodson:1948:SPI

[Dod48] Joseph C. Dodson. The statistical program of Iowa State College. The American Statistician, 2(3):13–14, June 1948. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682084.

Eckler:1948:PCP

[Eck48] A. Ross Eckler. Planning for the 1950 censuses of population and agriculture. The American Statistician, 2(1):6–10, February 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682360.

Eisenhart:1948:RSC

[Eis48] Churchill Eisenhart. The role of a statistical consultant in a research organization. The American Statistician, 2(2):6–7, April 1948. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682227.

Ekebald:1949:SSE

[Eke49] Frederick A. Ekebald. Should scales extend to zero? A case for the affirmative. The American Statistician, 3(1):24–25, February 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681395.

Epstein:1947:SAH

[Eps47] Albert Epstein. Statistical analysis with hand punched and sorted cards. The American Statistician, 1(2):6–7, October 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681622.

Givens:1949:NFS

[Giv49] Meredith B. Givens. Notes on federal–state relations in statistics. With illustrative reference to employment, unemployment and the labor market. The American Statistician, 3(5):6–9, December 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682407.

Guye:1948:PAP

[Guy48] Roberto Guye. A program of action for promoting statistical education in the Western Hemisphere. The American Statistician, 2(2): 10–13, April 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682229.

Haemer:1947:HLP

[Hae47a] Kenneth W. Haemer. Hold that line. A plea for the preservation of chart scale rulings. The American Statistician, 1(1):25, August 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681528.

Haemer:1947:PP

[Hae47b] Kenneth W. Haemer. The perils of perspective. The American Statistician, 1(3):19, December 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682537.

Haemer:1948:DSD

[Hae48a] Kenneth W. Haemer. Double scales are dangerous. The American Statistician, 2(3):24, June 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682088.

Haemer:1948:RBC

[Hae48b] Kenneth W. Haemer. Range-bar charts. The American Statistician, 2(2):23, April 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682234.

Haemer:1949:ABM

[Hae49a] Kenneth W. Haemer. Area bias in map presentation. The American Statistician, 3(2):19, April/May 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682742.

Haemer:1949:SSC

[Hae49b] Kenneth W. Haemer. The supplementary-scale chart. Two charts for the price of one. The American Statistician, 3(3):11, June/ July 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682140.

Harriman:1947:WIS

[Har47] W. Averell Harriman. A welcome to the International Statistical Conferences. The American Statistician, 1(1):4–5, August 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681519.

Hattery:1947:PPP

[Hat47] Lowell H. Hattery. Personnel placement by professional societies. The American Statistician, 1(2):10–13, October 1947. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681624.

Hayes:1948:ICD

[Hay48] Samuel P. Hayes, Jr. Industrial classification by Dun and Bradstreet. The American Statistician, 2(4):8–19, August 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682682.

Hegarty:1948:HUS

[Heg48] Edward J. Hegarty. How to use a set of display charts. The American Statistician, 2(5):32, October 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681651.

Hitt:1948:TFC

[Hit48] Homer L. Hitt. Three-factor cartographic representation. The American Statistician, 2(1):21–22, February 1948. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682365.

Howell:1948:SCA

[HJ48] J. M. Howell and Lee Johnson. Statistical control of accidents. The American Statistician, 2(1):16–19, February 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682363.

Hoadley:1947:CN

[Hoa47] Walter E. Hoadley, Jr. Chapter notes. The American Statistician, 1(1):20–22, August 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681525.

Horvath:1948:ORS

[Hor48] William J. Horvath. Operations research — a scientific basis for executive decisions. The American Statistician, 2(5):6–19, October 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681644.

Jahnel:1948:WUP

[Jah48] Frederick Jahnel. When to use pictorial symbols. The American Statistician, 2(4):25, August 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682691.

Jensen:1948:GSU

[Jen48] Vera Jensen. Graduate school, U.S. Department of Agriculture, sponsors internships in sampling. The American Statistician, 2 (4):20–21, August 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682687.

Kelley:1949:SCA

[Kel49] Truman Lee Kelley. Suiting the chart to the audience. Common graphic devices classified according to ease of reading. The American Statistician, 3(5):11, December 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682409.

Kendall:1948:WDL

[Ken48a] M. G. Kendall. Who discovered the Latin square? The American Statistician, 2(4):13, August 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682684.

Kendall:1948:ERS

[Ken48b] Maurice G. Kendall. The examination of the Royal Statistical Society. The American Statistician, 2(2):8–15, April 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682228.

Kyd:1948:DBS

[Kyd48] James Gray Kyd. Doomsday Book for Scotland. The American Statistician, 2(5):9–15, October 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681645.

Leavens:1949:OMS

[Lea49] Dickson H. Leavens. Overlapping membership of societies. The American Statistician, 3(3):4–5, June/July 1949. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682138.

Lehman:1948:SGE

[Leh48] John W. Lehman. Statistical gaps and economic indicators. The American Statistician, 2(5):10–12, October 1948. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681646.

Lieu:1948:CSC

[Lie48] D. K. Lieu. Collecting statistics in china. The American Statistician, 2(6):12–13, December 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681556.

Linnenberg:1949:DFSa

[Lin49a] Clem C. Linnenberg, Jr. The development of Federal Statistical Coordination, 1908–1949. The American Statistician, 3(2):6–11, April/May 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682737.

Linnenberg:1949:DFSb

[Lin49b] Clem C. Linnenberg, Jr. The development of Federal Statistical Coordination, 1908–1949. The American Statistician, 3(3):6–10, June/July 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682139.

Mosteller:1948:QAf

[MA48] Frederick Mosteller and R. L. Anderson. Questions and answers. The American Statistician, 2(6):16–17, December 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681558.

Mosteller:1948:QAc

[MBM^{+}48] Frederick Mosteller, G. W. Brown, A. M. Mood, Esdras Minville, Leo A. Aroian, and C. Horace Hamilton. Questions and answers. The American Statistician, 2(3):22–23, June 1948. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682087.

Mosteller:1949:QAb

[ME49] Frederick Mosteller and Lillian Epstein. Questions and answers. The American Statistician, 3(2):18, April/May 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682741.

Mosteller:1947:QAb

[MF47] Frederick Mosteller and Harold Freeman. Questions and answers. The American Statistician, 1(2):12–13, October 1947. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681625.

Mosteller:1948:QAe

[MF48] Frederick Mosteller and R. A. Fisher. Questions and answers: Answer to question 14 on combining independent tests of significance. The American Statistician, 2(5):30–31, October 1948. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681650.

Mosteller:1948:QAd

[MGG48] Frederick Mosteller, Roe Goodman, and E. J. Gumbel. Questions and answers. The American Statistician, 2(4):22–23, August 1948.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682689.

Mosteller:1948:QAb

[MH48] Frederick Mosteller and Kenneth W. Haemer. Questions and answers. The American Statistician, 2(2):18–19, April 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682232.

Mosteller:1949:QAe

[MMH49] Frederick Mosteller, Philip J. McCarthy, and Kenneth Haemer. Questions and answers. The American Statistician, 3(5):10, December 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682408.

Mosteller:1947:QAa

[Mos47] Frederick Mosteller. Questions and answers. The American Statistician, 1(1):22, August 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681526.

Mosteller:1949:QAc

[Mos49] Frederick Mosteller. Questions and answers. The American Statistician, 3(3):12–13, June/July 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682141.

Mosteller:1949:QAd

[MT49] Frederick Mosteller and John W. Tukey. Questions and answers. The American Statistician, 3(4):12, October 1949. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681357.

Mosteller:1948:QAa

[MWF48] Frederick Mosteller, Charles P. Winsor, and C. H. Fischer. Questions and answers. The American Statistician, 2(1):18–19, February 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682364.

Mosteller:1949:QAa

[MWK49] Frederick Mosteller, Emmett H. Welch, and V. S. Kolesnikoff. Questions and answers. The American Statistician, 3(1):22–23, February 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681393.

Myers:1947:LDL

[Mye47] R. J. Myers. Laymen don’t have logarithmic eyes. The case against the casual use of semi log charts. The American Statistician, 1(2): 17, October 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681628.

Neiswanger:1947:WRC

[NA47] William A. Neiswanger and H. K. Allen. A well-rounded curriculum in statistics. The American Statistician, 1(1):16–19, August 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681524.

Peach:1947:QA

[PTM47] Paul Peach, John W. Tukey, and Frederick Mosteller. Questions and answers. The American Statistician, 1(3):17–18, December 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682536.

Reed:1948:STS

[RCV48] Vergil D. Reed, Katherine G. Capt, and Herbert A. Vitriol. Selection, training, and supervision of field interviewers in marketing research. The American Statistician, 2(3):15–20, June 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682085.

Rice:1949:ISI

[Ric49] Stuart A. Rice. The International Statistical Institute. The American Statistician, 3(2):4–5, April/May 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682736.

Seaton:1948:SMM

[Sea48] G. L. Seaton. The statistician and modern management. The American Statistician, 2(6):10–11, December 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681555.

Shaw:1948:RSS

[Sha48] Ralph R. Shaw. Royal Society Scientific Information Conference. The American Statistician, 2(4):14–23, August 1948. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682685.

Stephan:1947:SOA

[SM47] Frederick F. Stephan and Philip J. McCarthy. Sampling opinion, attitudes, and consumer wants. The American Statistician, 1(3):6–7, December 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682532.

Snedecor:1948:DSI

[Sne48] George W. Snedecor. On the design of sampling investigations. The American Statistician, 2(6):6–13, December 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681554.

Society:1948:TSE

[Soc48] The Royal Statistical Society. The teaching of statistics in England. The American Statistician, 2(1):12–20, February 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682362.

Stebbins:1948:PSS

[Ste48a] Kathleen Stebbins. Placement service of the Special Libraries Association. The American Statistician, 2(4):21, August 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682688.

Sterns:1948:IWB

[Ste48b] F. H. Sterns. If I were the boss. The American Statistician, 2 (1):11–20, February 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682361.

Stevens:1948:PNY

[Ste48c] Milt Stevens. Perspective from New York. The American Statistician, 2(6):21, December 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681561.

Stevens:1949:PNYa

[Ste49a] Milt Stevens. Perspective from New York. The American Statistician, 3(1):23, February 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681394.

Stevens:1949:PNYb

[Ste49b] Milt Stevens. Perspective from New York. The American Statistician, 3(2):11, April/May 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682738.

Taeuber:1948:FPS

[Tae48] Conrad Taeuber. FAO’s program for statistical training. The American Statistician, 2(6):14–18, December 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681557.

Tukey:1947:QIO

[Tuk47] John W. Tukey. Question 4: Infinite observations. The American Statistician, 1(3):17–18, ???? 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tukey:1949:AQM

[Tuk49a] John W. Tukey. Answer to question 21: Maximum likelihood estimates, efficiency. The American Statistician, 3(4):12, ???? 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Tukey:1949:MSFa

[Tuk49b] John W. Tukey. Memorandum on statistics in the Federal Government. Part I — Chapters I–IV. The American Statistician, 3(1):6–17, February 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681391.

Tukey:1949:MSFb

[Tuk49c] John W. Tukey. Memorandum on statistics in the Federal Government. Part II, Chapters V–VII. The American Statistician, 3(2):12–16, April/May 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682739.

Tumbleson:1947:SE

[Tum47] Robert C. Tumbleson. Statistical engineering. The American Statistician, 1(1):7–11, August 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681521.

Uri:1948:GRM

[Uri48] George Uri. Graphic representation for multiple linear correlation. The American Statistician, 2(6):19, December 1948. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681559.

Watkins:1947:CN

[Wat47] J. H. Watkins. Chapter notes. The American Statistician, 1(2):14, October 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681626.

Wiers:1948:CC

[Wie48] Paul Wiers. Criminal comparisons. The American Statistician, 2(3): 11–12, June 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682083.

Wilks:1947:CSP

[Wil47a] S. S. Wilks. Crisis in statistical personnel. The American Statistician, 1(2):8–16, October 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681623.

Willcox:1947:LSS

[Wil47b] Walter F. Willcox. Lemuel Shattuck, statist founder of the American Statistical Association. The American Statistician, 1(1):11–13, August 1947. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681522.

Yocum:1949:LE

[YA49] James C. Yocum and Sam Arnold. Letter to the Editor. The American Statistician, 3(3):1, June/July 1949. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682136.

Zapoleon:1948:OWS

[Zap48] Marguerite Wykoff Zapoleon. The outlook for women in statistics. The American Statistician, 2(5):16–29, October 1948. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2681648.